OpenStudy (anonymous):

Find the value of z: -x+y+2z=20 x-4y+4z=13 -2x-6y-3z= -10

OpenStudy (anonymous):

add and subtract equations.

OpenStudy (anonymous):

@wio teach me how

OpenStudy (anonymous):

you use 3 var elimination

OpenStudy (anonymous):

add the first and the second equation and what do you get ?

OpenStudy (anonymous):

yeah .. @mth3v4 is right..

OpenStudy (anonymous):

Make it a matrix\[ \begin{bmatrix} -1&1&2 &20\\\ 1&-4&4&13\\ -2&-6&-3&-10 \end{bmatrix} \]

OpenStudy (anonymous):

\[ R_1+R_2\to R_2 \]

OpenStudy (anonymous):

Make it a matrix\[ \begin{bmatrix} -1&1&2 &20\\\ 0&-3&6&33\\ -2&-6&-3&-10 \end{bmatrix} \]

OpenStudy (anonymous):

\[ -2R_1+R_3\to R_3 \]

OpenStudy (anonymous):

\[ \begin{bmatrix} -1&1&2 &20\\\ 0&-3&6&33\\ 0&-8&-9&-50 \end{bmatrix} \]

OpenStudy (anonymous):

You keep going like this.

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

-x + y + 2z = 20 x - 4y + 4z = 13 ---------------add 0 - 3y + 6z = 33 -3y + 6z = 33 now add the 2nd and the third equation. But since we cancelled out the x in the first two equations, we want to make the x's cancel out these equations. x - 4y + 4y = 13-->(2) -2x - 6y - 3z = -10 ------------------- 2x - 8y + 8y = 26 (result of multiplying by 2) -2x - 6y - 3z = -10 -------------------add 0 - 14y + 5z = 16 -14y + 5z = 16 now take the answers you got and try to eliminate a variable -3y + 6z = 33 -->(5) -14y + 5z = 16 -->(-6) -------------------- -15y + 30z = 165 (result of multiplying by 5) 84y - 30z = - 96 (result of multiplying by -6) --------------------add 69y = 69 y = 1 now sub 1 in for y into either of the equations that does not have a x in it -14y + 5z = 16 -14(1) + 5z = 16 -14 + 5z = 16 5z = 16 + 14 5z = 30 z = 6 now sub 1 in for y and 6 in for z into any of the original equations -x + y + 2z = 20 -x + 1 + 2(6) = 20 -x + 1 + 12 = 20 -x + 13 = 20 -x = 20 - 13 -x = 7 x = -7 check... x - 4y + 4z = 13 -7 - 4(1) + 4(6) = 13 -7 - 4 + 24 = 13 -11 + 24 = 13 13 = 13 (correct) x = -7, y = 1, z = 6

OpenStudy (anonymous):

wio is probably doing it the easier way by using the matrix.

OpenStudy (anonymous):

however which way you still are going to perform the same more likely operations ?